Circular No. 6238

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J2058+42 C. A. Wilson, S. N. Zhang, M. H. Finger, R. B. Wilson, and M. Scott,

Marshall Space Flight Center, NASA; and T. Koh, D. Chakrabarty, B. Vaughan, and T. A. Prince, California Institute of Technology, report for the Compton Observatory BATSE team: “Pulsations from a previouslyuncatalogued source in the Cygnus region in the energy band 20–50 keV have been detected since Sept. 14 and continue to be detected through Sept.

25. A preliminary location (95-percent confidence) is bounded by a quadrilateral with vertices (equinox 2000.0) at α = 21^{h}03^{m}, δ = +40^{o}.2; 20^{h}56^{m}, +44^{o}.0; 20^{h}53^{m}, +43^{o}.2; 21^{h}01^{m}, +39^{o}.4. Positions estimated independently from the pulsed and total flux are consistent. An rms pulsed flux of 130 ± 40 mCrab in the energy band 20–50 keV was observed on Sept. 23. Observations between Sept. 20 and 23 show pulsations up to at least 70 keV and are consistent with a barycentric pulse frequency of 5.05618(7) × 10^{−3} Hz on Sept. 22.0 UT with a frequency derivative of 1.70(7) × 10^{−11} Hz/s. The total flux, measured by Earth occultation, increased from < 100 mCrab on Sept. 17 to 280 ± 50 mCrab on Sept. 22 in the band 20–100 keV and has remained approximately constant (within 15 percent) from Sept. 22 to 24. The energy spectrum of the total flux on Sept. 20–24 was well fit by an optically-thin thermal bremsstrahlung model with kT = 17 ± 1 keV from 20 to 100 keV.”

SUPERNOVA 1995ac IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports the following pre-

cise position for SN 1995ac: α = 22^{h}45^{m}34^{s}.14, δ = −8^{o}45′04′′.7 (equinox 2000.0), which is 2′′.1 west and 4′′.5 south of the galaxy’s center. A nearby star (V ∼ 16.5) has position end figures 38^{s}.41, 44′27′′.3.

COMET C/1995 Q1 (BRADFIELD) Total visual magnitude estimates: Aug. 26.37 UT, 5.2 (R. H. Mc-

Naught, Coonabarabran, N.S.W., 10×50 binoculars); Sept. 23.52, 7.6 (C. S. Morris, southern CA, 0.26-m reflector); 25.18, 7.2 (J. D. Shanklin, Cambridge, England, 14×100 binoculars); 26.45, 7.7 (G. W. Kronk, Troy, IL, 0.33-m reflector); 27.16, 6.5 (G. A. Milani, Mt. Grappa, Italy, 20×80 binoculars); 27.51, 7.7 (Morris, 20×80 binoculars).

1995 September 27 Daniel W. E. Green